OpenStudy (anonymous):

Jake is comparing the prices of two laundry service companies. Company A charges $20 per load and an additional $8 as service charges. Company B charges $24 per load and an additional $4 as service charges. Part A: Write equations to represent Company A's and Company B's total laundry charges for a certain number of loads. Define the variable used in the equations. Part B: Which company would charge less for 5 loads of laundry? Justify your answer. Part C: How much money is saved by using the services of Company A instead of Company B to clean 9 loads of laundry?

OpenStudy (flvs.net):

Company A: 20x + 8 Company B: 24x + 4 where x = # of loads Part A: see above Part B: For 5 loads, Company A would be20(5) + 8 = $108 Company B would be 24(5) + 4 = $124 Company A would charge less. PART C: For 9 loads, Company A: 20(9) + 8 = $188 Company B: 24(9) + 4 = $220 You would save $32 by using Company A.

OpenStudy (aaronandyson):

Company A $20 per load $8 service Company B $24 per load $4 service A) Let l be the number of loads Company A = 20 l + 8 Company B = 24 l + 4 B) Company A = 20*5 + 8 = $108 Company B = 24*5 + 4 = $124 Company A charges less C) Company A = 20*9 + 8 = $188 Company B = 24*9 + 4 = $220 220 - 188 = 32 You save $32.

OpenStudy (flvs.net):

Good job @AaronAndyson

OpenStudy (anonymous):

thnx guys :)

OpenStudy (flvs.net):

Np

OpenStudy (aaronandyson):

YW!